# Robust Radiotherapy Planning with Spatially Based Uncertainty Sets

### v1

@@ 5.3 Dose-Volume Experiments
In this section, we conduct experiments using Algorithm 2, introduced in Section 4 for handling dose-volume constraints. The experiments in this segments were run using Julia 1.3.0, and Gurobi solver 9.0.0, on an Intel Xeon Gold 6254 CPU with a 3.10GHz base frequency, 24.75 MB cache, and 376 RAM.
In order to speed up the computation, we may first narrow down the search interval further by performing a preliminary bisection search, and then proceed by running Algorithm 5 given this reduced interval. Note that invoking the bisection search may sacrifice some of the theoretical properties guaranteed by the initial lower bound, thus serving only as an effective heuristic. Further, in order to deal with possible degeneracy in Algorithm 2 we determine adjacency of basic feasible solutions only based on the $x$ and $y$ variables. In these experiments, we focus on the dose-volume constraint for the brain (OAR 2 of Patient 4). Specifically, following the ICRU guidelines (ICR 2010), $99\%$ of the brain should receive less than $\bar{d}_{2}=60$ Gy while $\hat{d}_{2}=62$ Gy; that is, $\alpha=0.01$ in constraint (11f). For these experiments we consider both the nominal and spatially robust settings with the same parameters as in Section 5.1.Recall that Algorithm 5 determines the minimal $\beta$ value for which the solution satisfies this constraint. Intuitively, a small $\beta$ and corresponding penalty term in the objective should correspond to a small change in the objective compared with the non-penalized problem (where $\beta=0$). Thus, the constraint is expected to be satisfied with a slight decline in the optimal objective value relative to that achieved when $\beta=0$.
Figure 12 depicts the value of the adjusted dose $\underaccent{\bar}{d}$, and the $\ell_{1}$ and $\ell_{0}$ norms of the deviation vector $y$, as a function of $\beta$, for the different values of $\beta$ obtained by the algorithm for both cases. In order to run the algorithm, we first computed $\bar{\beta}\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.22}\times 10^{-2}$ for the nominal case, and $\bar{\beta}\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.58}\times 10^{-2}$ for the specially robust case. We used $\underaccent{\bar}{\beta}=0$ in both cases, while the true lower bounds are $\underaccent{\bar}{\beta}\approx 1.73\times 10^{-6}$ for the nominal setting and $\underaccent{\bar}{\beta}\approx 1.36\times 10^{-6}$ for the robust setting. The algorithm determined that $\beta^{*}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.72}\times 10^{-5}$ and $\beta^{*}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2.4}\times 10^{-5}$ are the smallest $\beta$ values that satisfy $\|y\|_{0}\leq\lfloor\alpha|H_{2}|\rfloor=13,601$ for the nominal and the spatially robust case, respectively (thus satisfying the dose-volume constraint).
@@
We compared the solutions obtain by Algorithm 5 in both of the spatially robust and nominal settings, with a solution obtained from problem (2.2) with an additional CVAR constraint, as suggested by Romeijn et al. (2003). Table 4 shows these results, comparing also with the unresricted case, where the dose-volume constraints are not imposed (having $\beta=0$ in (13)).
@@ 6 Conclusions and Future Work
We proposed a novel spatially based uncertainty set to model uncertainty in tumor radiosensitivity of radiation treatment based on measured biomarkers. Naively incorporating this uncertainty set in robust voxel-based FMO models would lead to an unmanageable increase in the already large problem size. Alternatively, we compactly reformulate the robust problem by proving that at most a few uncertainty set extreme points need to be considered for each constraint. While theoretically tractable, the new compact formulation may still contain over a billion constraints. We first show that although the compact formulation cannot be solved directly it can be solved within a reasonable running time by carefully generating constraints as needed. Extensive experiments demonstrate advantages of the proposed uncertainty set in terms of the adjusted homogeneity-dose tradeoff over the nominal and robust box-uncertain models.
@@
This work has focused on spatial aspects of biomarker related uncertainty in a voxel-level planning model (also known as “dose painting”). In future work, we would like to extend the proposed model to include temporal aspects, taking into account fractionation of treatments as well as a possible additional image scan midtreatment. This leads to a challenging two-stage adaptive robust problem, where the spatial uncertainty changes over time. Finally, it is worth noting that the study of robust and adaptive planning under the condition that neighboring elements uncertainty is connected has application beyond radiation therapy. Communicable disease, traffic congestion, and even agriculture are all examples of situations where uncertainty among neighbors in space and time is connected.
@@
All data used in this paper is based on data available from public repositories whose references are indicated and with processing as described in Section 5. This data in its final processed form will be made available by the authors upon reasonable request.
@@ A.2 Row and Column Generation Algorithm
Here, $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ denotes an instance of problem (13) with constraints (10d) and (10e) defined only for a subset $S\subseteq T^{2}$, constraints (11c) defined only for subsets $\hat{H}_{1}\subseteq H_{1},\ldots,\hat{H}_{K-1}\subseteq H_{K-1}$, and constraints (11d),(11e) defined for subset $\hat{H}_{K}\subseteq H_{K}$.
@@ B.2 Penalty Parameter Bounds Lemma
Proof.  For the following proofs, we use notation $\mathcal{X}$ to denote the feasible solution set of (13). Also, for every $\Theta\geq 0$, $Z_{C}(\Theta)$ denotes the optimal value of problem (4) with the right-hand side of constraint (12c) given by $\Theta$. Finally, consider the dual function of (4), obtained by dualizing constraint (12c) with dual multiplier $\beta$,
@@ Appendix C Conversion of PET Image SUV to OMF
In our experiments, imaging standard uptake values (SUV) are read directly from an FMISO-PET scans. The SUV values are typically normalized with respect to well oxygenated region, and in our case the reference value of $pO2=26$ is fixed as the the median over the entire brain (excluding the tumor) [McKeown, 2014]. These normalized SUV readings are converted to oxygen pressure $pO2$ values using the formula
@@
where $A=10.9$, $B=10.7$ and $C=2.5$ [Toma-Dasu et al., 2012, Lindblom et al., 2017]. Then, the OER is given by
@@
see for example Titz and Jeraj [2008]. The values assumed here are $K=3$ and $m=3$ following Saka et al. [2014]. Finally, the OMF is simply OER normalized as a proportion of the maximum OER value in the image. Finally, we apply interpolation to deduce voxel OMF values for the CT images that may have a higher resolution of voxels than the corresponding PET images.
Note that the values of the constants in this appendix appear to be highly uncertain according to the scientific literature. In particular, for the conversion of $pO2$ to $OER$Titz and Jeraj [2008] consider a range of possible values $K=2,3,4$. Note that for $pO2=15$ it implies that the corresponding $OMF$ value could vary in the interval $(0.875,0.930)$ (assuming a well oxygenated reference point has $pO2=151$) due only to parameter $K^{\prime}s$ uncertainty, while significant uncertainty may be associated with the values of other parameters as well. Based on such a sensitivity analysis, experimenting with values of $\delta$ in $[0,0.14]$ would be reasonable.
@@ Appendix D Additional Numerical Results
Figure 13: Optimal solution value of (2.2.2) for $\gamma=0.04$ and different values of $\delta$.
Figure 14: Nominal performance of optimal solutions to (2.2.2) with different values of $\gamma$ and $\delta$, compared with the nominal solution.
Figure 15: Physical performance of optimal solutions to (2.2.2) with different values of $\gamma$ and $\delta$, compared with the nominal solution.
Figure 16: Worst-case performance ($\hat{\mu}$, $\hat{\underline{d}}$) of solutions optimal to (2.2.2) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.02$ and $\delta$ as indicated. The performance of the nominal solution is also shown for comparison.
Figure 17: Worst-case performance ($\hat{\mu}$, $\hat{\underline{d}}$) of solutions optimal to (2.2.2) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.04$ and $\delta$ as indicated. The performance of the nominal solution is also shown for comparison.

### v2

@@ 5.3 Dose-Volume Experiments
In this section, we conduct experiments using Algorithm 2, introduced in Section 4 for handling dose-volume constraints. rThe experiments in this segments were run using Julia 1.3.0, and Gurobi solver 9.0.0, on an Intel Xeon Gold 6254 CPU with a 3.10GHz base frequency, 24.75 MB cache, and 376 RAM.
In order to speed up the computation, we may first narrow down the search interval further by performing a preliminary bisection search, and then proceed by running Algorithm 5 given this reduced interval. Note that invoking the bisection search may sacrifice some of the theoretical properties guaranteed by the initial lower bound, thus serving only as an effective heuristic. Further, in order to deal with possible degeneracy in Algorithm 2 we determine adjacency of basic feasible solutions only based on the $x$ and $y$ variables. In these experiments, we focus on the dose-volume constraint for the brain (OAR 2 of Patient 4). Specifically, following the ICRU guidelines (ICR, 2010), $99\%$ of the brain should receive less than $\bar{d}_{2}=60$ Gy while $\hat{d}_{2}=62$ Gy; that is, $\alpha=0.01$ in constraint (11f). For these experiments we consider both the nominal and spatially robust settings with the same parameters as in Section 5.1.Recall that Algorithm 5 determines the minimal $\beta$ value for which the solution satisfies this constraint. Intuitively, a small $\beta$ and corresponding penalty term in the objective should correspond to a small change in the objective compared with the non-penalized problem (where $\beta=0$). Thus, the constraint is expected to be satisfied with a slight decline in the optimal objective value relative to that achieved when $\beta=0$.
Figure 12 depicts the value of the adjusted dose $\underaccent{\bar}{d}$, and the $\ell_{1}$ and $\ell_{0}$ norms of the deviation vector $y$, as a function of $\beta$, for the different values of $\beta$ obtained by the algorithm for both cases. In order to run the algorithm, we first computed $\bar{\beta}\approx{2.22}\times 10^{-2}$ for the nominal case, and $\bar{\beta}\approx{2.58}\times 10^{-2}$ for the specially robust case. We used $\underaccent{\bar}{\beta}=0$ in both cases, while the true lower bounds are $\underaccent{\bar}{\beta}\approx 1.73\times 10^{-6}$ for the nominal setting and $\underaccent{\bar}{\beta}\approx 1.36\times 10^{-6}$ for the robust setting. The algorithm determined that $\beta^{*}=r{2.72}\times 10^{-5}$ and $\beta^{*}={2.4}\times 10^{-5}$ are the smallest $\beta$ values that satisfy $\|y\|_{0}\leq\lfloor\alpha|H_{2}|\rfloor=13,601$ for the nominal and the spatially robust case, respectively (thus satisfying the dose-volume constraint).
@@
We compared the solutions obtain by Algorithm 5 in both of the spatially robust and nominal settings, with a solution obtained from problem (3) with an additional CVAR constraint, as suggested by Romeijn et al., (2003). Table 4 shows these results, comparing also with the unresricted case, where the dose-volume constraints are not imposed (having $\beta=0$ in (13)).
@@ 6 Conclusions and Future Work
We proposed a novel spatially based uncertainty set to model uncertainty in tumor radiosensitivity of radiation treatment based on measured biomarkers. Naively incorporating this uncertainty set in robust voxel-based FMO models would lead to an unmanageable increase in the already large problem size. Alternatively, we compactly reformulate the robust problem by proving that at most a few uncertainty set extreme points need to be considered for each constraint. While theoretically tractable, the new compact formulation may still contain over a billion constraints. We first show that although the compact formulation cannot be solved directly, it can be solved within a reasonable running time and consumption of memory, by carefully generating constraints as needed. Extensive experiments demonstrate advantages of the proposed uncertainty set in terms of the adjusted homogeneity-dose tradeoff over the nominal and robust box-uncertain models.
In this work, we also develop a method for satisfying a single dose-volume constraint that outperforms the standard CVaR modeling approach to this problem. Although handling dose volume constraints remains computationally challenging for clinical use, this can be addressed by future research focused on voxel aggregation techniques for solving large scale instances within reasonable running times.
@@
This work has focused on spatial aspects of biomarker related uncertainty in a voxel-level planning model (also known as “dose painting”). In future work, we would like to extend the proposed model to include temporal aspects, taking into account fractionation of treatments as well as a possible additional image scan midtreatment. This leads to a challenging two-stage adaptive robust problem, where the spatial uncertainty changes over time. Finally, it is worth noting that the study of robust and adaptive planning, under the condition that the uncertainty of neighboring elements is connected, has application beyond radiation therapy. Communicable disease, traffic congestion, and even agriculture are all examples of situations where uncertainty among neighbors in space and time is connected.
Acknowledgments
The authors are grateful to Aditya P. Apte for his help in using the CERR software to process and export the imaging datasets used in this paper. Noam Goldberg acknowledges the computational resources provided by the Data Science Institute at Bar-Ilan University.
Notes on the Contributors
Noam Goldberg
is a senior lecturer in the Department of Management at Bar-Ilan University and an incoming associate professor at the Department of Industrial Engineering and Management at Ben-Gurion University. He completed a PhD in operations research at Rutgers University (RUTCOR), New Brunswick, NJ, in 2010. His graduate studies followed an early non-academic system and software engineering career in telecommunications. His PhD research on sparse data classification models was awarded a first prize in a contest of the INFORMS NJ chapter in 2009. He held several postdoctoral positions during 2009-2013 at the Technion-Israel Institute of Technology, Argonne National Laboratory and Carnegie Mellon University. He joined Bar-Ilan University in 2013 as a tenure-track lecturer, where he became a tenured senior lecturer since 2018. In fall 2019 he was a visiting associate professor at the Rutgers Business School. He has more than 30 peer reviewed publications including top journals in optimization and conference proceedings in computer science. In current work he continues to be motivated by high-impact applications of OR, data driven optimization, game models and solution methods.
Mark Langer, MD
is Professor of Clinical Radiation Oncology at Indiana University School of Medicine, Indianapolis, IN USA. His research interests lie in mathematical programming and optimization for radiation treatment planning and design, and clinical strategies to improve outcomes in head and neck cancers. He is the recipient of grants from the NIH, NSF and the Indiana 21 st Century Fund and has served on review panels for the NIH, NSF, BSF and the DOD. He served as chair for the Operations Research and Radiotherapy Collaborative Working Group (ORART) sponsored by the NIH and was a subawardee on the ORART Collaborative Working Group’s Tool Kit program funded by the NSF. He has authored publications in Ann. Operations Research, Mathematical Programming B, Linear Algebra and its Applications, J. Operat. Res. Soc., Journal of Mechanisms and Robotics, Journal of Applied Clinical Medical Physics, Dysphagia, Int. J. Radiat. Oncol. Biol. Phys., Radiotherapy and Oncology, IIE Transactions on Health Care Systems Engineering, Phys. Med. Biol, and Medical Physics.
Shimrit Shtern
is a senior lecturer in the Faculty of Data and Decision Sciences at the Technion - Israel Institute of Technology. She completed a PhD in optimization at the Technion in 2015. She was post-doctoral fellow at the Operation Research Center at MIT in 2015-2017 and joined the Technion as a faculty member in 2018. Her research focuses on theory and applications of optimization under uncertainty as well as development and convergence analysis of first order methods for structured optimization problems. Her work has been published in top operation research and optimization journals including Management Science, Operations Research, Mathematics of Operations Research, SIAM Journal on Optimization, and Mathematical Programming.
@@ Acknowledgments
All data used in this paper is based on data available from public repositories whose references are indicated and with processing as described in Section 5. The processed data and code are available at https://github.com/ShimShtern/RobustRTPStatic.
@@ A.2 Row and Column Generation Algorithm
Here, $\Phi(S,\hat{H}_{1},\ldots,\hat{H}_{K},\beta)$ denotes an instance of problem (13) with constraints (10c) and (10d) defined only for a subset $S\subseteq T^{2}$, constraints (11c) defined only for subsets $\hat{H}_{1}\subseteq H_{1},\ldots,\hat{H}_{K-1}\subseteq H_{K-1}$, and constraints (11d),(11e) defined for subset $\hat{H}_{K}\subseteq H_{K}$.
@@ B.2 Penalty Parameter Bounds Lemma
For the following proofs, we use notation $\mathcal{X}$ to denote the feasible solution set of (13). Also, for every $\Theta\geq 0$, $Z_{C}(\Theta)$ denotes the optimal value of problem (12) with the right-hand side of constraint (12c) given by $\Theta$. Finally, consider the dual function of (12), obtained by dualizing constraint (12c) with dual multiplier $\beta$,
@@ Appendix C Conversion of PET Image SUV to OMF
In our experiments, imaging standard uptake values (SUV) are read directly from an FMISO-PET scans. The SUV values are typically normalized with respect to well oxygenated region, and in our case the reference value of $pO2=26$ is fixed as the the median over the entire brain (excluding the tumor) (McKeown, 2014). These normalized SUV readings are converted to oxygen pressure $pO2$ values using the formula
@@
where $A=10.9$, $B=10.7$ and $C=2.5$ (Toma-Dasu et al., 2012 and Lindblom et al., 2017). Then, the OER is given by
@@
see for example Titz and Jeraj, (2008). The values assumed here are $K=3$ and $m=3$ following Saka et al., (2014). Finally, the OMF is simply OER normalized as a proportion of the maximum OER value in the image. Finally, we apply interpolation to deduce voxel OMF values for the CT images that may have a higher resolution of voxels than the corresponding PET images.
Note that the values of the constants in this appendix appear to be highly uncertain according to the scientific literature. In particular, for the conversion of $pO2$ to $OER$Titz and Jeraj, (2008) consider a range of possible values $K=2,3,4$. Note that for $pO2=15$ it implies that the corresponding $OMF$ value could vary in the interval $(0.875,0.930)$ (assuming a well oxygenated reference point has $pO2=151$) due only to parameter $K^{\prime}s$ uncertainty, while significant uncertainty may be associated with the values of other parameters as well. Based on such a sensitivity analysis, experimenting with values of $\delta$ in $[0,0.14]$ would be reasonable.
@@ Appendix D Additional Numerical Results
Figure 13: Optimal solution value of (10) for $\gamma=0.04$ and different values of $\delta$.
Figure 14: Nominal performance of optimal solutions to (10) with different values of $\gamma$ and $\delta$, compared with the nominal solution.
Figure 15: Physical performance of optimal solutions to (10) with different values of $\gamma$ and $\delta$, compared with the nominal solution.
Figure 16: Worst-case performance ($\hat{\mu}$, $\hat{\underline{d}}$) of solutions optimal to (10) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.02$ and $\delta$ as indicated. The performance of the nominal solution is also shown for comparison.
Figure 17: Worst-case performance ($\hat{\mu}$, $\hat{\underline{d}}$) of solutions optimal to (10) with misspecified $\gamma$, under the assumption that the “true” $\mathcal{U}_{\text{SB}}$ has $\gamma=0.04$ and $\delta$ as indicated. The performance of the nominal solution is also shown for comparison.